OpenStudy (anonymous):

The fibonacci sequence is defined recurively by the following a1=1 a2=1 an=an-1-an-2 A.true b.false

hartnn (hartnn):

next term is not the difference, its the sum., the correct one is : an=an−1+an−2

OpenStudy (anonymous):

so its false

hartnn (hartnn):

yep, false

OpenStudy (anonymous):

let bn=bn-1,b1=1Then b7 equals A.28 B.7 C.42 D.7+6

hartnn (hartnn):

\(b_n\) is what ?

OpenStudy (anonymous):

on the n+b the n-1 should be slightly below the b but idn how to type it like that

hartnn (hartnn):

is \(\Large b_n = n+ b_{n-1}\) ?

OpenStudy (anonymous):

yes,sorry

hartnn (hartnn):

we know b1 =1 find b2 plug in n=2 in that equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

whats the answer?

hartnn (hartnn):

could you find b2 ? then find b3, b4...and so on till b7

OpenStudy (anonymous):

couldnt find it

hartnn (hartnn):

what have you tried ?

OpenStudy (anonymous):

is it bn=2+1?

hartnn (hartnn):

b2 =2+1 = 3 yes now find b3 b3 =3 + b2 = 3+3 = ... similarly, find b4,b5,..

OpenStudy (anonymous):

would the answer be A

OpenStudy (anonymous):

wait ..... wouldnt it be b7=6+1=7?

hartnn (hartnn):

no

hartnn (hartnn):

\(\Large b_7 = 7+b_6 = 7+21=28\)

OpenStudy (anonymous):

thats what i said the first time b7=28

OpenStudy (wolf1728):

Fibonacci sequence (as others have said) progresses as 1 1 2 3 5 8 13 21 34 55 http://1728.org/fibonacci.htm

hartnn (hartnn):

and thats correct :)

OpenStudy (anonymous):

whats the first 5 tems of the fibonacci sequence generated by the fibonacci sequence rule below a1=3 a2=-2 a3=an-1+an-2 A.3,-2,1,-1,0 B.-2,3,1,4,5 C.3,-2,1,3,4 D.3,-2,5,3,8

OpenStudy (wolf1728):

I have no idea. I know about the Fibonacci Sequence but not when they start adding new rules to it.

OpenStudy (anonymous):

i cant figure it out

OpenStudy (anonymous):

i have no idea